# GRAPES: Learning to Sample Graphs for Scalable Graph Neural Networks

### v2

@@ Appendix B Experimental Details
For all experiments, we used as architecture the Graph Convolutional Network [23], with two layers, a hidden size of 256, batch size of 256, and a sampling size of 256 nodes per layer. We implemented the GCNs in GRAPES via PyTorch Geometric [12]. We train for 50 epochs on Cora, Citeseer, and Reddit; 100 epochs on BlogCat, DBLP, Flickr, ogbn-products, Pubmed, snap-patents, and Yelp; and 150 epochs on ogbn-arxiv and ogbn-proteins. The ogbn-proteins and BlogCat datasets do not contain node features, and instead we learn node embeddings for them of dimension 128 for ogbn-proteins, and 64 for BlogCat.
@@ B.1 Hyperparameter Tuning
We tune the hyperparameters of GRAPES using a random search strategy with the goal of maximizing the accuracy of the validation dataset. We used Weights and Biases for hyperparameter tuning 333 https://wandb.ai . The best-performing hyperparameters for every dataset can be found in our repository https://github.com/dfdazac/grapes. The following are the hyperparameters that we tuned: the learning rate of the GFlowNet, the learning rate of the classification GCN, and the scaling parameter $\alpha$. We used the log uniform distribution to sample the aforementioned hyperparameters with the values from the following ranges respectively, $[1e-6,1e-2]$, $[1e-6,1e-2]$, and [1e2, 1e6]. We kept the other hyperparameters such as the batch size and hidden dimension of the GCN. We used the Adam optimizer [22] for $\operatorname{GCN}_{\operatorname{C}}$ and $\operatorname{GCN}_{\operatorname{S}}$.
@@ B.2 Baselines
FastGCN [9] LADIES [51] GraphSAINT [47] AS-GCN [18] GAS [13]
@@
For LADIES we used the official implementation, which also contains an implementation of FastGCN. We changed the nonlinear activation function from ELU to ReLU, and we removed any linear layers after the two layers of the GCN, set dropout to zero, and disabled early stopping. We also noticed that the original LADIES implementation divided the target nodes into mini-batches, not from the entire graphs as we do, but into random fragments. This means that LADIES and FastGCN do not see all the target nodes in the training data. We kept this setting unchanged because otherwise it significantly slowed down the training of these two methods.
For GraphSAINT we noticed that the GNN consists of two layers of higher order aggregator, which is a combination of GraphSage-mean [16] and MixHop [2], and a linear classification layer at the end. Moreover, the original implementation of GraphSAINT is only applicable to inductive learning on the graphs, where the train graph only contains the train nodes and is entirely different from the validation and test graphs where only the nodes from the validation set and test are available respectively. We argue that in transductive learning, unlike inductive learning, the motivation to scale to larger graphs is higher since the validation and test nodes are also available during training, and therefore, the processed graph is larger. Finally, to keep all the configurations the same as GRAPES, we used PyTorch Geometric’s function for GraphSAINT Random Walk sampler with depth two, which showed the best performance across the three variations of GraphSAINT. Therefore, we use the same GCN and data loader (transductive) as ours and use GraphSAINT to sample a subgraph of nodes for training. We use the node sampler setting since it is the only setting that allows specifying different sampling budgets, and therefore, can be compared to layer-wise methods with the same sampling budget. We also removed the early stopping and use the same number of epochs as GRAPES. Note that for Figure 3, we need to change the sample size while Random Walk sampler can only sample equal to the batch size per layer. Therefore, for these experiments we used the Node Sampler version of GraphSAINT, which allows us to change the sample size. For instance, for sample size $32$ per layer, we set GraphSAINT’s Node Sampler’s sample size to $256+32+32=320$ to sample a subgraph with 320 nodes. Please refer to our repository for more details about the implementation of GraphSAINT.
@@ Appendix C Dataset statistics
We present the statistics of the datasets used in our experiments in Table 4. The splits that we used for Cora, Citeseer, and Pubmed correspond to the “full” splits, in which the label rate is higher than in the “public” splits. For BlogCat, we take the average accuracy of all the methods across the three available splits provided by [48]. For DBLP and snap-patents we use the average of ten random splits because these two datasets had no predefined splits.
Table 4: Statistics of the datasets used in our experiments. The label rate indicates the percentage of nodes used for training. ogbn-proteins and BlogCat do not contain node features, and instead we learn embeddings for nodes in these datasets.
@@ Appendix D Entropy as Node preference measure
Figures 5 and 6 show the mean and standard deviation of entropy in base two of all the datasets. We calculate the mean entropy for as the following:
@@
Figure 5: Combined entropy plots for Citeseer, Cora, Flickr, ogbn-arxiv, ogbn-products, Pubmed, Reddit and Yelp showcasing the mean entropy. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
Figure 6: Combined entropy plots for BlogCat, snap-patents, DBLP and ogbn-proteins, showcasing the mean entropy across epochs. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
@@ Appendix E GPU Memory usage comparison between GRAPES and GAS
We compared different variants of GRAPES, with different sample sizes (32, 256), with GAS [13], which is a non-sampling method. Figure 2 shows the GPU memory allocation (MB), in a logarithmic scale for GRAPES-32, GRAPES-256 and GAS. The three graph methods exhibit distinct performance characteristics across various datasets. We used the max_memory_allocated function in PyTorch to measure the GPU memory allocation. [^4] https://pytorch.org/docs/stable/generated/torch.cuda.max_memory_allocated.html Since this function measures the maximum memory allocation since the beginning of the program, where the memory measurement is done is not relevant.
@@ F.1 Generative Flow Networks
Generative Flow Networks (GFlowNets) [6, 7] are generative models that can sample from a very large structured space. GFlowNets construct the structure in multiple generation steps. Compared to Reinforcement Learning approaches, GFlowNets learn to sample in proportion to a given reward function, while in Reinforcement Learning, reward functions are maximized. This feature of GFlowNets encourages sampling diverse sets of high-quality structures, instead of only considering the single best structure. GFlowNets have been used in several applications like molecule design and material science [6, 14, 20], Bayesian structure learning [11], scientific discovery [21], and GNN explainability [26]. Similar to our work, the latter utilizes a GFlowNet to sample subgraphs. However, this method explains a trained GNN and is not used to scale GNN training to large graphs.
@@ F.2 GFlowNet and Trajectory Balance Loss
We first give a brief overview of GFlowNets [6, 7] and the trajectory balance loss [28]. Let $\mathcal{G}_{F}=(\mathcal{S},\mathcal{A},\mathcal{S}_{0},\mathcal{S}_{f},R)$ denote a GFlowNet learning problem. Here, $\mathcal{S}$ is a finite set of states that forms a directed graph with $\mathcal{A}$, a set of directed edges representing actions or transitions between states. $\mathcal{S}_{0}\subset\mathcal{S}$ is the set of initial states, $\mathcal{S}_{f}\subset\mathcal{S}$ is the set of terminating states, [^5] Technically, GFlowNets have unique source and terminal (or ‘sink’) states $s_{s}$ and $s_{f}$. The source state has an edge to all initial states, and all terminating states have an edge to the terminal state. and $R:\mathcal{S}_{f}\rightarrow\mathbb{R}_{+}$ is the reward function defined on terminating states. At time $t$, a particular $a_{t}\in\mathcal{A}$ indicates the action taken to transition from state $s^{(t-1)}$ to $s^{(t)}$. A trajectory $\tau$ is a path through the graph from an initial state $s^{(0)}$ to a terminating state $s^{(n)}\in\mathcal{S}_{f}$: $\tau=(s^{(0)}\to\ldots\to s^{(n)})$. A GFlowNet is a neural network that learns to transition from an initial state $s^{(0)}$ to a terminating state where the reward $R(s^{(n)})$ is given. The goal of the GFlowNet is to ensure that following the forward transition probabilities $P_{F}(s^{(t)}|s^{(t-1)})$ leads to final states $s\in\mathcal{S}_{f}$ with probability in proportion to the reward $R$ [6]. The Trajectory Balance (TB) loss [28] is developed with this goal. For a trajectory $\tau=(s^{(0)}\to\ldots\to s^{(n)})$, the TB loss is:
@@
where $Z:\mathcal{S}_{0}\rightarrow\mathbb{R}_{+}$ computes the total flow of the network from the starting state $s^{(0)}$ and $P_{F}$ and $P_{B}$ are the forward and the backward transition probabilities between the states, where both can be parameterized by a neural network [28].
@@ F.4 Reward scaling
In our experiments, we noticed that with the bigger datasets, the GFlowNet is more affected by the log-probabilities than the reward from the classification $\operatorname{GCN}_{\operatorname{C}}$. The reason that the term $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ dominates $\mathcal{L}_{\operatorname{GFN}}$ is that $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})=\sum_{i\in\mathcal{N}(K^{(l-1)})}\log p_{i}$, which sums over $|\mathcal{N}(K^{(l-1)})|$ elements. Given a batch size of $256$, this neighborhood can be as big as $52\hskip 1.49994pt000$ nodes, resulting in summing $52\hskip 1.49994pt000$ log-probabilities. The majority of the probabilities $p_{i}$ are values close to zero. Therefore, the above sum would be a large negative number. Since the loss $\mathcal{L}_{\operatorname{C}}$ is, in our experiments, often quite close to zero, the log-probability and its variance dominate the loss. Therefore, we add the hyper-parameter $\alpha$ to the reward and tune it in our experiments.

### v3

@@ Appendix B Experimental Details
For all experiments, we used as architecture the Graph Convolutional Network (Kipf & Welling, 2016), with two layers, a hidden size of 256, a batch size of 256, and a sampling size of 256 nodes per layer. We implemented the GCNs in GRAPES via PyTorch Geometric (Fey & Lenssen, 2019). We train for 50 epochs on Cora, Citeseer, and Reddit; 100 epochs on BlogCat, DBLP, Flickr, ogbn-products, Pubmed, snap-patents, and Yelp; and 150 epochs on ogbn-arxiv and ogbn-proteins. The ogbn-proteins and BlogCat datasets do not contain node features, and instead we learn node embeddings for them of dimension 128 for ogbn-proteins, and 64 for BlogCat.
@@ B.1 Hyperparameter Tuning
We tune the hyperparameters of GRAPES using a random search strategy with the goal of maximizing the accuracy of the validation dataset. We used Weights and Biases for hyperparameter tuning 555 https://wandb.ai . The best-performing hyperparameters for every dataset can be found in our repository https://anonymous.4open.science/r/GRAPES. The following are the hyperparameters that we tuned: the learning rate of the GFlowNet, the learning rate of the classification GCN, and the scaling parameter $\alpha$. We used the log uniform distribution to sample the aforementioned hyperparameters with the values from the following ranges, respectively, $[1e-6,1e-2]$, $[1e-6,1e-2]$, and [1e2, 1e6]. We kept the other hyperparameters, such as the batch size and hidden dimension of the GCN. We used the Adam optimizer (Kingma & Ba, 2014) for $\operatorname{GCN}_{\operatorname{C}}$ and $\operatorname{GCN}_{\operatorname{S}}$.
We did a hyperparameter sensitivity analysis performed using Weights & Biases Biewald (2020) for training of GRAPES-GFN on Yelp. Importance measures how strongly each parameter influences the validation accuracy (higher = more critical), while correlation shows the direction and magnitude of the linear relationship (positive = increasing the parameter tends to increase the accuracy, negative = the opposite). From the table below, $lr_{gc}$, the learning rate for the classification GCN, has the strongest effect on accuracy and is negatively correlated, suggesting that higher learning rates can harm performance. Meanwhile, $lr_{gf}$ and $\alpha$ show smaller impacts and weaker correlations, indicating that the model is less affected by changes in these two parameters.
Table 3: Importance and correlation of configuration parameters for training GRAPES-GFN on Yelp.
Config Parameter Importance Correlation
lr_gc 0.633 -0.610
lr_gf GCN) 0.280 0.055
$\alpha$ 0.087 0.097
@@ B.2 Baselines
FastGCN (Chen et al., 2018b) LADIES (Zou et al., 2019) GraphSAINT (Zeng et al., 2019) AS-GCN (Huang et al., 2018) GAS (Fey et al., 2021)
@@
For LADIES, we used the official implementation, which also contains an implementation of FastGCN. We changed the nonlinear activation function from ELU to ReLU, and we removed any linear layers after the two layers of the GCN, set dropout to zero, and disabled early stopping. We also noticed that the original LADIES implementation divided the target nodes into mini-batches, not from the entire graphs as we do, but into random fragments. This means that LADIES and FastGCN do not see all the target nodes in the training data. We kept this setting unchanged because otherwise it would have significantly slowed down the training of these two methods.
For GraphSAINT, we noticed that the GNN consists of two layers of higher-order aggregators, which are a combination of GraphSage-mean (Hamilton et al., 2017) and MixHop (Abu-El-Haija et al., 2019), and a linear classification layer at the end. Moreover, the original implementation of GraphSAINT is only applicable to inductive learning on the graphs, where the training graph only contains the training nodes and is entirely different from the validation and test graphs, where only the nodes from the validation set and test are available, respectively. We argue that in transductive learning, unlike inductive learning, the motivation to scale to larger graphs is higher since the validation and test nodes are also available during training, and therefore, the processed graph is larger. Finally, to keep all the configurations the same as GRAPES, we used PyTorch Geometric’s function for GraphSAINT Random Walk sampler with depth two, which showed the best performance across the three variations of GraphSAINT. Therefore, we use the same GCN and data loader (transductive) as ours and use GraphSAINT to sample a subgraph of nodes for training. We use the node sampler setting since it is the only setting that allows specifying different sampling budgets, and therefore, can be compared to layer-wise methods with the same sampling budget. We also removed the early stopping and used the same number of epochs as GRAPES. Note that for Figure 4, we need to change the sample size while the Random Walk sampler can only sample equal to the batch size per layer. Therefore, for these experiments, we used the Node Sampler version of GraphSAINT, which allows us to change the sample size. For instance, for sample size $32$ per layer, we set GraphSAINT’s Node Sampler’s sample size to $256+32+32=320$ to sample a subgraph with 320 nodes. Please refer to our repository for more details about the implementation of GraphSAINT.
@@
For PASS, we changed the inference to full-batch to match the rest of the baselines. We set the parameter sampling_scope to $512$, because in the original experiments of the paper, this value was set to twice the batch size. For Flickr, ogbn-arxiv, and ogbn-proteins, this value results in OOM; therefore, we set it to $128$, $128$, and $256$, respectively. For the datasets that we report OOM in tables 1 and 2, we tested all possible sampling_scope between $512$ and $16$.
@@ Appendix C Dataset statistics
We present the statistics of the datasets used in our experiments in Table 5. The splits that we used for Cora, Citeseer, and Pubmed correspond to the “full” splits, in which the label rate is higher than in the “public” splits. For BlogCat, we take the average accuracy of all the methods across the three available splits provided by (Zhao et al., 2023). For DBLP and snap-patents, we use the average of ten random splits because these two datasets had no predefined splits.
Table 5: Statistics of the datasets used in our experiments. The label rate indicates the percentage of nodes used for training. ogbn-proteins and BlogCat do not contain node features, and instead, we learn embeddings for nodes in these datasets.
@@ Appendix D Entropy as Node preference measure
Figures 6 and 7 show the mean and standard deviation of entropy in base two of all the datasets. We calculate the mean entropy as the following:
@@
Figure 6: Combined entropy plots for Citeseer, Cora, Flickr, ogbn-arxiv, ogbn-products, Pubmed, Reddit, and Yelp showcasing the mean entropy. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
Figure 7: Combined entropy plots for BlogCat, snap-patents, DBLP, and ogbn-proteins, showcasing the mean entropy across epochs. The shaded region indicates the standard deviation of the entropy across nodes. The plots compare 1-hop node sampling against 2-hop node sampling.
We analyzed the label distribution of the sampled subgraphs via GRAPES-GF for Yelp, by counting the percentage of nodes having each label and calculating the difference between the original and sampled graphs, i.e., percentage of nodes having label $c$ in the original graph minus percentage of nodes having label $c$ in the sampled graph. As shown in the figure, all the values are positive, indicating that a higher ratio of nodes have each label in the original graph. Moreover, this difference increases at the end of training. This means that among the 100 labels in Yelp, GRAPES prefers the nodes with fewer labels or that are even single-labeled. This may make the classification task easier and result in a lower loss. Moreover, among the 100 classes, some are sampled the most and some the least. Figure 8 below shows this trend.
Figure 8: Comparison of Label distributions between GRSAPES-GFN and the original graph for Yelp. The left figure shows the percentage of nodes in the last epoch of both graphs that have each label. The right shows the difference between those percentages in the original graph and the sampled graph at the beginning and end of training.
@@ Appendix E GPU Memory usage comparison between GRAPES and GAS
We compared different variants of GRAPES, with different sample sizes (32, 256), with GAS (Fey et al., 2021), which is a non-sampling method. Figure 3 shows the GPU memory allocation (MB), on a logarithmic scale for GRAPES-32, GRAPES-256, and GAS. The three graph methods exhibit distinct performance characteristics across various datasets. We used the max_memory_allocated function in PyTorch to measure the GPU memory allocation. [^6] https://pytorch.org/docs/stable/generated/torch.cuda.max_memory_allocated.html Since this function measures the maximum memory allocation since the beginning of the program, where the memory measurement is done is not relevant.
@@ F.1 Generative Flow Networks
Generative Flow Networks (GFlowNets) (Bengio et al., 2021a; b) are generative models that can sample from a very large structured space. GFlowNets construct the structure in multiple generation steps. Compared to Reinforcement Learning approaches, GFlowNets learn to sample in proportion to a given reward function, while in Reinforcement Learning, reward functions are maximized. This feature of GFlowNets encourages sampling diverse sets of high-quality structures, instead of only considering the single best structure. GFlowNets have been used in several applications like molecule design and material science (Bengio et al., 2021a; Gao et al., 2022; Jain et al., 2022), Bayesian structure learning (Deleu et al., 2022), scientific discovery (Jain et al., 2023), and GNN explainability (Li et al., 2023). Similar to our work, the latter utilizes a GFlowNet to sample subgraphs. However, this method explains a trained GNN and is not used to scale GNN training to large graphs.
@@ F.2 GFlowNet and Trajectory Balance Loss
We first give a brief overview of GFlowNets (Bengio et al., 2021a; b) and the trajectory balance loss (Malkin et al., 2022a). Let $\mathcal{G}_{F}=(\mathcal{S},\mathcal{A},\mathcal{S}_{0},\mathcal{S}_{f},R)$ denote a GFlowNet learning problem. Here, $\mathcal{S}$ is a finite set of states that forms a directed graph with $\mathcal{A}$, a set of directed edges representing actions or transitions between states. $\mathcal{S}_{0}\subset\mathcal{S}$ is the set of initial states, $\mathcal{S}_{f}\subset\mathcal{S}$ is the set of terminating states, [^7] Technically, GFlowNets have unique source and terminal (or ‘sink’) states $s_{s}$ and $s_{f}$. The source state has an edge to all initial states, and all terminating states have an edge to the terminal state. and $R:\mathcal{S}_{f}\rightarrow\mathbb{R}_{+}$ is the reward function defined on terminating states. At time $t$, a particular $a_{t}\in\mathcal{A}$ indicates the action taken to transition from state $s^{(t-1)}$ to $s^{(t)}$. A trajectory $\tau$ is a path through the graph from an initial state $s^{(0)}$ to a terminating state $s^{(n)}\in\mathcal{S}_{f}$: $\tau=(s^{(0)}\to\ldots\to s^{(n)})$. A GFlowNet is a neural network that learns to transition from an initial state $s^{(0)}$ to a terminating state where the reward $R(s^{(n)})$ is given. The goal of the GFlowNet is to ensure that following the forward transition probabilities $P_{F}(s^{(t)}|s^{(t-1)})$ leads to final states $s\in\mathcal{S}_{f}$ with probability in proportion to the reward $R$ (Bengio et al., 2021a). The Trajectory Balance (TB) loss (Malkin et al., 2022a) is developed with this goal. For a trajectory $\tau=(s^{(0)}\to\ldots\to s^{(n)})$, the TB loss is:
@@
where $Z:\mathcal{S}_{0}\rightarrow\mathbb{R}_{+}$ computes the total flow of the network from the starting state $s^{(0)}$ and $P_{F}$ and $P_{B}$ are the forward and the backward transition probabilities between the states, where both can be parameterized by a neural network (Malkin et al., 2022a).
@@ F.4 Reward scaling
In our experiments, we noticed that with the bigger datasets, the GFlowNet is more affected by the log-probabilities than the reward from the classification $\operatorname{GCN}_{\operatorname{C}}$. The reason that the term $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$ dominates $\mathcal{L}_{\operatorname{GFN}}$ is that $\log q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})=\sum_{i\in\mathcal{N}(K^{(l-1)})}\log p_{i}$, which sums over $|\mathcal{N}(K^{(l-1)})|$ elements. Given a batch size of $256$, this neighborhood can be as big as $52,000$ nodes, resulting in summing $52,000$ log-probabilities. The majority of the probabilities $p_{i}$ are values close to zero. Therefore, the above sum would be a large negative number. Since the loss $\mathcal{L}_{\operatorname{C}}$ is, in our experiments, often quite close to zero, the log-probability and its variance dominate the loss. Therefore, we add the hyperparameter $\alpha$ to the reward and tune it in our experiments.
Appendix G GRAPES and the straight-through estimator
Table 6: Comparison between GRAPES and variants using the straight-through estimator (STE).
Cora Flickr ogbn-arxiv Yelp
Best GRAPES 87.62 $\pm$ 0.48 49.54 $\pm$ 0.67 62.58 $\pm$ 0.64 44.57 $\pm$ 0.88
STE 87.95 $\pm$ 0.18 46.33 $\pm$ 1.39 61.95 $\pm$ 0.32 37.29 $\pm$ 0.15
[…]
The straight-through estimator Bengio et al. (2013) (STE) has been proposed as a method for learning in computation graphs involving a non-differentiable function $f$ of an input $x$, by replacing $f$ with the identity in the backward pass during computation of the gradients.
In practice, applying the STE for learning in GRAPES requires ensuring that the gradients flow through the sampled nodes, for example by treating the top-k sampled nodes at each layer as a mask which multiplies the messages from these nodes to all their neighbors. Such a computation over all neighboring nodes defeats the purpose of sampling, increasing memory usage.
[…]
Appendix H Proof of Theorem 1
H.1 Proof
In this section, we prove Theorem 1 in section 5 as follows:
Proof.
Let $\mathcal{G}=(\mathcal{V,E)}$ be an undirected fully connected simple graph with an even number of nodes N, $\mathcal{V}=\{1,\dots,N\}$, a set of edges =$\mathcal{E}=\{e_{ij}\}_{i,j=1}^{N}$, where $e_{ij}$ denotes an edge between $v_{i}$ and $v_{j}$. Let $Y=\{y_{i}|i\in\{1,\dots,N\},y_{i}\in\{0,1\}\}$ be the set of node labels and let $X\in\mathbb{R}^{N\times 2}$ be the node features. Let $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ be one partition of $\mathcal{E}$, where $\mathcal{E}_{1}$ consists of $N/2$ edges with two conditions: 1) $e_{ij}\in\mathcal{E}_{1}$, if and only if the first feature of nodes $v_{i}$ and $v_{j}$ are equal, i.e. $e_{ij}\in\mathcal{E}_{1}\Leftrightarrow x_{i1}=x_{j1}$, and 2) each node appears only at one edge in $\mathcal{E}_{1}$. The second condition implies that each value for the first feature appears only in the nodes connected by one edge in $\mathcal{E}_{1}$, i.e., for two distinct edges $e_{ij},e_{kl}\in\mathcal{E}_{1}$, $x_{1i}\neq x_{1k}$.
For each edge $e_{ij}$ in $\mathcal{E}_{1}$, sample $x_{i2}\sim\text{Bern}(0.5)$ on $\{0,1\}$. Then, set the node label $y_{j}$ to $x_{i2}$. That is, the label of each node is equal to the second feature of its corresponding neighbor in $\mathcal{E}_{1}$. Since $\mathcal{G}$ is fully connected, the $L$-hop neighborhood for all nodes is $\mathcal{V}$. Therefore, non-adaptive methods cannot prefer any node over another: a GCN with a non-adaptive sampler can only achieve accuracy above chance by sampling $K$ nodes such that for each target node, the neighbor with the same first feature is sampled. The probability of sampling such a neighbor is $p=\frac{\binom{N-2}{K-1}}{\binom{N-1}{K}}=\frac{K}{N-1}$. The nominator indicates choosing the desired neighbor connected to the target node in $\mathcal{E}_{1}$ and then choosing the remaining $K-1$ nodes among the $N-2$ nodes. The denominator indicates choosing $K$ nodes among $N-1$ nodes (excluding the target node). Repeating the sampling for $L$ layers, eliminating $K$ nodes from the candidate neighbors for sampling for the next layer, will result in $p=\sum_{i=0}^{L-1}{\frac{K}{N-iK-1}}\leq\frac{LK}{N}$.
However, there is a GCN with an adaptive sampler that classifies nodes with an accuracy of $100\%$. Specifically, such an adaptive sampler compares the first node feature of the neighbors with the target node $v_{i}$, and then always samples the unique $v_{j}$ where $e_{ij}\in\mathcal{E}_{1}$. Subsequently, it samples $K-1$ additional neighbors randomly. If the sampler perfectly samples the informative neighbors in $\mathcal{E}_{1}$, we assume there exists a GCN that can classify the target nodes correctly. We show a construction of such a GCN in section H.2. Figure 9 shows an example of such a graph with eight nodes and a homophily ratio of $0.43$. ∎
Figure 9: An example of a graph for Theorem 1 with eight nodes. Red edges belong to $\mathcal{E}_{1}$, features $x_{i}$ and labels $y_{i}$ are shown beside every node. For nodes $v_{1}$ and $v_{2}$ we show the edge $e_{12}$ as an example. As shown, the label of each node is the second feature of its neighbor, where a red edge connects them. The edge homophily ratio is $h=\frac{12}{28}=0.43$.
In the above proof, we show that there are graphs for which there is a performance gap between adaptive and non-adaptive samplers. Usually $LK\ll N$; therefore, the probability $p\leq\frac{LK}{N}$ for non-adaptive sampling methods to perform better than the chance level becomes very small. These graphs are heterophilous, with an expected homophily ratio equal to $0.5$. We calculate this expectation below:
$\mathbb{E}(h_{\mathcal{G}})=\mathbb{E}\left(\frac{\binom{m}{2}+\binom{N-m}{2}}{\binom{N}{2}}\right),m\sim\text{Binomial}(N,0.5)$
For that, we need $\mathbb{E}\left(\binom{m}{2}\right)=\frac{1}{2}\mathbb{E}\left(m(m-1)\right)$. Because $\mathbb{E}(m)=\frac{N}{2}$ and $Var(m)=\frac{N}{4}$, we have
$\mathbb{E}(m(m-1))=\mathbb{E}(m^{2}-m)=\mathbb{E}(m^{2})-\mathbb{E}(m)=Var(m)+\left(\mathbb{E}(m)\right)^{2}-\mathbb{E}(m)$
$=\frac{N}{4}+\left(\frac{N}{2}\right)^{2}-\frac{N}{2}=\frac{N^{2}-N}{4},$
so we have
$\mathbb{E}\left(\binom{m}{2}\right)=\frac{1}{2}\times\frac{N^{2}-N}{4}=\frac{N^{2}-N}{8}.$
Similarly, $\mathbb{E}\left(\binom{N-m}{2}\right)=\frac{N^{2}-N}{8}$. Therefore, the expected homophily ratio is
$\mathbb{E}(h_{\mathcal{G}})=\frac{\frac{N^{2}-N}{8}+\frac{N^{2}-N}{8}}{\binom{N}{2}}=0.5.$
H.2 GCN Construction
In this section, we construct a GCN that can compute the classification rule assumed in the proof of Theorem 1. Let $\mathbf{x}_{i}\in\mathbb{R}^{2}$ be an input feature vector of a node $v_{i}\in\mathcal{V}$. We construct a GCN where the first layer computes a 4-dimensional vector for a target node $v_{i}$:
$\mathbf{h}_{i}^{(1)}=\tilde{\mathbf{W}}_{1}\mathbf{x}_{i}+\frac{1}{|\mathcal{N}(v_{i})|}\sum_{v_{j}\in\mathcal{N}(v_{i})}\mathbf{W}_{1}\mathbf{x}_{j}.$ (13)
Here, $\tilde{\mathbf{W}}_{1}$ is the self-loop weight, and $\mathbf{W}_{1}$ is the weight for the neighbors of a target node. We define the following:
$\tilde{\mathbf{W}}_{1}=\begin{bmatrix}1&0\\ 0&1\\ 0&0\\ 0&0\end{bmatrix},\quad\mathbf{W}_{1}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&1\end{bmatrix}$ (14)
For a given target node, the output of this layer is a vector $\mathbf{h}_{i}^{(1)}\in\mathbb{R}^{4}$ where the first two values correspond to the target node’s features, and the other two values are the average of the features of its neighbors. The next layer implements the following function: if values $h_{i1}^{(1)}$ and $h_{i3}^{(1)}$ are the same, then the output is the second value of this vector, otherwise it is zero. We denote this function as $h_{i}^{(2)}=f(\mathbf{h}_{i}^{(1)})\in\mathbb{R}$, which can be implemented with a multi-layer perceptron due to its universal approximation properties (Hornik et al., 1989). An MLP can be implemented in a GCN by omitting the neighbors and using a self-loop weight matrix only. The last layer of the GCN selects the average value in the neighborhood of a target node from the scalars computed by $f$, which can be computed using a simple GCN where the self-loop weight is set to zero: $h_{i}^{(3)}=\frac{1}{|\mathcal{N}(v_{i})|}\sum_{v_{j}\in\mathcal{N}(v_{i})}h_{j}^{(2)}$. This output corresponds to the classification rule required for Theorem 1.